# Slab tilings, flips and the triple twist

### v1

@@ 5 On connected components
Figure 20: A tiling which admits no flips. We have $\operatorname{Tw}_{\kappa}=+4$ for $\kappa$ equal to red-green.
Let us quickly review the situation for domino tilings. We are interested in the connected components under flips of the space of domino tilings of a given box. There are often connected components of size 1, i.e., tilings which admit no flip. Figure 1 shows an example for the $3\times 3\times 2$ box; there are similar examples in larger boxes.
@@
Figure 21: Another tiling which admits no flips. In this case, $\operatorname{Tw}_{\kappa}=+12$.
@@
There is a constant $C_{1}>0$ such that the following assertion holds: If ${\cal R}=[0,N]^{3}$ is a cube with even side lengths (and thus slab tileable) then $\lvert\operatorname{TTw}[{\cal R}]\rvert\leq C_{1}N^{12}$.
@@
Given a pair of colors $\kappa$, we bound $\lvert\operatorname{Tw}_{\kappa}[{\cal R}]\rvert$. The transformed region $\tilde{{\cal R}}$ is contained in a cube of side $N$ (see Figures 18 and 22 for an example of the transformed region for $N=4$). Fix $f$ a floor of $\tilde{{\cal R}}$, fix a column $c$ of $f$, and use axes as in Figure 22. Consider dominoes with a square in $c$: there are $d_{y}$ dominoes in the direction of the $y$-axis and $d_{z}$ dominoes in the direction of the $z$-axis. The number of possible effects happening within $c$ is at most $d_{y}d_{z}$. Since $d_{y}+d_{z}\leq N$, we have $d_{y}d_{z}\leq\frac{N^{2}}{4}$ by AM-GM inequality.
$y$
$x$
Figure 22: A typical floor after transforming a $4\times 4\times N$ box. It is bounded by the dotted $4\times 4$ square. We show $3$ dominoes on the central column. In this example, $d_{y}=d_{z}=1$ and there is one effect.
Now, varying $c$ and then $f$, we see there are at most $\frac{N^{4}}{4}$ effects within $\tilde{{\cal R}}$. Thus, for every tiling ${\mathbf{t}}$ we have $|\operatorname{Tw}_{\kappa}({\mathbf{t}})|\leq\frac{N^{4}}{16}$. Since $N\geq 2$, we have
This completes the proof, with $C_{1}=\frac{1}{2^{8}}$. ∎
Lemma 5.4.
Proof.
Consider the subdivision of the $[0,6n]\times[0,6n]\times[0,8n]$ box in parts $\alpha,\beta$ and $\gamma$ as in Figure 23. We tile $\alpha$ with horizontal slabs. We can tile $\beta$ and $\gamma$ both as collections of $2n^{2}$ annuli as in Figure 24.
$y$
$x$
$\alpha$
$\alpha$
$\beta$
$\alpha$
$\alpha$
$\alpha$
$\alpha$
$\beta$
$\beta$
$\gamma$
$\alpha$
$\alpha$
$\alpha$
$\alpha$
$\alpha$
$\gamma$
$\gamma$
$\beta$
$\alpha$
$\alpha$
$\gamma$
Figure 23: Floor diagram of subdivision of the $[0,6n]\times[0,6n]\times[0,8n]$ box. There are four kinds of floors. Each of them appears $2n$ times. Each rectangle on each floor has dimensions either $6n\times 2n$, $2n\times 6n$ or $2n\times 2n$.
Consider the surfaces $S_{\beta}$ and $S_{\gamma}$ which are both squares of side $2n$, depicted as red cuts in Figure 24. Look at $S_{\beta}$ from the $y$-axis view. We regard each floor of $S_{\beta}$ with only grey pieces (a grey floor) as a domino tiling of an annulus, and we consider the flux with respect to the red cut. The flux across $S_{\beta}$ is the sum of the fluxes on all grey floors. The flux across $S_{\gamma}$ is defined analogously. We see the fluxes across each of the surfaces $S_{\beta}$ and $S_{\gamma}$ are integers in the interval $[-n^{2},n^{2}]$. We can choose the fluxes across $S_{\beta}$ and $S_{\gamma}$ to be equal to $u$ and $v$, respectively. We call this tiling ${\mathbf{t}}={\mathbf{t}}_{u,v}$.
$S_{\gamma}$
$z$
$x$
$S_{\beta}$
$y$
$z$
Figure 24: Example of $\beta$ ($y$-axis view) and $\gamma$ ($x$-axis view) on the $[0,12]\times[0,12]\times[0,16]$ box. Note that the top left vertex of a floor on the upper (resp. lower) diagram is of the form $(0,*,0)$ (resp. $(*,0,0)$). The surfaces $S_{\beta}$ and $S_{\gamma}$ are depicted in red.
We now prove $\operatorname{Tw}_{\kappa_{z}}({\mathbf{t}})=\pm 2uv$. We will consider $\tilde{{\mathbf{t}}}$ as a dimer covering ${\mathbf{t}}^{*}$ on the graph ${\cal R}^{*}$ whose vertices are centers of squares in $\tilde{{\cal R}}$ and edges join centers of adjacent squares. We draw ${\cal R}^{*}$ on top of ${\cal R}$ for ease of visualization. We put red dots on the squares of ${\cal R}$ painted with good colors. We proceed to define surfaces $\tilde{S}_{\beta}$ and $\tilde{S}_{\gamma}$ as follows. From the $z$-axis view, look at the first floor of $\beta$ which intersects $S_{\beta}$. Consider the red dots immediately above or below the red cut and draw a polygonal line joining them from left to right as in Figure 25. Then, $\tilde{S}_{\beta}$ is a curtain with base this polygonal line and height $12n-$ 1. The definition of the surface $\tilde{S}_{\gamma}$ is analogous: we now consider red dots to the left or right of the red cut, also in Figure 25.
$x$
$y$
Figure 25: Construction of the surfaces $\tilde{S}_{\beta}$ (above) and $\tilde{S}_{\gamma}$ (below) following the example in Figure 24. Both are $z$-views. The polygonal lines of $\tilde{S}_{\beta}$ and $\tilde{S}_{\gamma}$ are dotted and the surfaces $S_{\beta}$ and $S_{\gamma}$ are represented as full lines.
To prove this, without loss, let ${\mathbf{t}}_{0}$ be ${\mathbf{t}}$ after twisting one of the annuli of $\beta$. The difference $\tilde{\mathbf{t}}_{0}-\tilde{{\mathbf{t}}}$ has one non trivial cycle which bounds an oriented surface $\tilde{S}$. We draw $\tilde{\mathbf{t}}_{0}-\tilde{{\mathbf{t}}}$ on top of ${\cal R}$ as in Figure 26. Notice $\tilde{S}$ is tangent to $\tilde{S}_{\gamma}$. To calculate $t=\operatorname{Tw}(\tilde{\mathbf{t}}_{0})-\operatorname{Tw}(\tilde{{\mathbf{t}}})$, we must count dominoes from $\tilde{{\mathbf{t}}}$ passing through $S$. These are exactly the ones which are part of an annulus in $\tilde{\gamma}$. By twisting one annulus from $\tilde{\gamma}$ we see $t$ changes by ±$\pm 2$, hence $t=\pm 2v$ by induction.
Figure 26: Some floors of the $[0,12]\times[0,12]\times[0,16]$ box. The surface $\tilde{S}$ bounded by $\tilde{\mathbf{t}}_{0}-\tilde{{\mathbf{t}}}$ is drawn as black polygonal lines. Full lines represent the first and last floors of this cycle. The first and last floors of two annuli from $\tilde{\gamma}$ are shown as purple polygonal lines. Full segments represent the actual dimers in the tiling.
$S_{\gamma}$
$S_{\beta}$
Figure 27: We continue with the example of Figure 24. On the first row we make the slabs in $\beta$ horizontal. On the second row we make the slabs in $\gamma$ and its hole horizontal.
Without loss of generality we treat the first case of the above claim. Refer to Figure 27. If this holds, then we can flip the slabs in $\beta$ so that all of them are horizontal. In particular, the slabs passing through $\gamma$’s hole are horizontal. Look at $\gamma$ from the $x$-axis view. For every pair of consecutive floors of $\gamma$ we can make all slabs horizontal by using the domino flip connectedness result for planar regions [4, 7]. ∎
Lemma 5.5.
Let ${\cal R}=[0,16n]^{3}$ be a box (which is slab tileable). We construct a slab tiling ${\mathbf{t}}$ of ${\cal R}$ such that $\operatorname{TTw}({\mathbf{t}})$ is any triple $t\in([-2n^{4},2n^{4}]\cap 2{\mathbb{Z}})^{3}$.
Proof.
For numbers $u_{i,j},v_{i,j}\in[-n^{2},n^{2}]$, $1\leq i\leq 2$, $1\leq j\leq 3$ we can construct tilings ${\mathbf{t}}_{i,j}$ of the $[0,6n]\times[0,6n]\times[0,8n]$ box such that $\operatorname{TTw}({\mathbf{t}}_{j})=2u_{i,j}v_{i,j}e_{j}$ by Lemma 5.4 (here $e_{j}\in{\mathbb{Z}}^{3}$ is the $j$-th vector in the canonical base). After doing this, we divide the $[0,16n]^{3}$ box in $8$ cubes of side $8n$. Each of these smaller cubes is tiled using one of the ${\mathbf{t}}_{i,j}$ on one corner, and horizontal slabs to fill the rest of the space. Two of the smaller cubes will receive only horizontal slabs. This new tiling is named ${\mathbf{t}}$.
Let $t=(t_{1},t_{2},t_{3})$. Choose
Here, for $s\in{\mathbb{R}}$, we have $s=\lfloor s\rfloor+\{s\}$ where $\lfloor s\rfloor\in{\mathbb{Z}}$ and $\{s\}\in[0,1)$. We have $t_{j}=2u_{1,j}v_{1,j}+2u_{2,j}v_{2,j}$, as desired. Then, for each $j$, $\operatorname{TTw}({\mathbf{t}}_{1,j}+{\mathbf{t}}_{2,j})=t_{j}e_{j}\in([-2n^{4},2n^{4}]\cap 2{\mathbb{Z}})^{3}$, where ${\mathbf{t}}_{1,j}+{\mathbf{t}}_{2,j}$ is the tiling of the two cubes containing ${\mathbf{t}}_{1,j}$ and ${\mathbf{t}}_{2,j}$. Summing the above for all $1\leq j\leq 3$ yields $\operatorname{TTw}({\mathbf{t}})=(t_{1},t_{2},t_{3})$. ∎
Lemma 5.6.
Proof.
Write $N=16M+2k$ for some $0\leq k\leq 7$. We can divide ${\cal R}$ into two parts ${\cal R}_{1}$ and ${\cal R}_{2}$ as in Figure 28. For a given $v\in([-2M^{4},2M^{4}]\cap 2{\mathbb{Z}})^{3}$, we tile ${\cal R}_{1}$ with a tiling ${\mathbf{t}}_{1}$ having $\operatorname{TTw}({\mathbf{t}}_{1})=v$ and we tile ${\cal R}_{2}$ with horizontal slabs.
$\mathcal{R}_{1}$
$\mathcal{R}_{2}$
Figure 28: Diagram of subdivision of ${\cal R}$. There is one kind of floor. It comprises one $16M\times 16M$ square on the lower left corner, ${\cal R}_{1}$, and its complement, ${\cal R}_{2}$.
Then, $\lvert\operatorname{TTw}[{\cal R}]\rvert\geq(2M^{4})^{3}=8M^{12}=8(\frac{N-2k}{16})^{12}=c(N-2k)^{12}$, where $c=\frac{1}{2^{45}}$. We seek $C>0$ such that $(N-2k)^{12}\geq CN^{12}$ holds. This happens if and only if $C\leq(1-\frac{2k}{N})^{12}$. By minimizing the RHS with the constraints $N\geq 256$ and $k\leq 7$, we can take $C=\frac{1}{2}\leq(1-\frac{14}{256})^{12}$. Finally, we can take $C_{0}=cC=\frac{1}{2^{46}}$. ∎
We summarize the previous lemmas in the proposition below.
Proposition 5.7.
If $N\geq 256$ is an even integer, then
Proof.
We take $C_{0}$ and $C_{1}$ as in the proofs of Lemmas 5.3 and 5.6. ∎
The flip is a natural local move. For domino tilings, there is another important local move besides the flip: the trit ([3]). The trit allows us to move between flip connected components. It is conjectured that any two tilings of a box can be connected by a finite sequence of flips and trits. It is known for large classes of boxes that this holds for almost all tilings [5]. It is also known that any tiling of a box admits at least one flip or trit [2]. In the case of slab tilings, however, there is no finite set of local moves with similar properties.
Proposition 5.8.
Consider the region ${\cal R}=[0,2k]\times[0,2k]\times[0,5l]$, $k\geq 3$, $l\geq 1$ and a slab tiling ${\mathbf{t}}$ following the pattern in Figure 20. Now consider ${\cal R}_{0}=[0,2(k-1)]\times[0,2(k-1)]\times[0,5l]$. In ${\cal R}$, there is only one way of removing and placing back the slabs which are contained in the interior of ${\cal R}_{0}$.
Proof.
Refer to Figure 29. We proceed by induction on $k$. We show the inductive process for five consecutive floors. The slab numbered $1$ is clearly stuck since it is “sandwiched” between two other slabs. Using this argument, on the first three floors, we proceed to deduce the positions of slabs which intersect the columns (in the floor diagram) between the dotted and red lines. In the figure, this corresponds to the slabs from $1$ to $8$. We repeat this, but now for slabs intersecting the rows between the dotted and red lines. This correspond to slabs from $9$ to $14$.
@@
Figure 29: We again use the tiling from Figure 21. The region ${\cal R}_{0}$ is bounded by the red curve. The region ${\cal R}_{0}=[0,2(k-2)]\times[0,2(k-2)]\times[0,5l]$ is bounded by the dotted curve.

### v2

@@ 6 Connectivity under local moves
Figure 20: A tiling which admits no flips; $\operatorname{Tw}_{\kappa}=+4$ for $\kappa$ equal to red-green
Let us recall a fact about domino tilings and flips: there are often connected components of size 1, i.e., tilings which admit no flip. Figure 1 shows an example for the $3\times 3\times 2$ box; there are similar examples in larger boxes (see [6] for examples).
@@
Figure 21: Another tiling which admits no flips; in this case, $\operatorname{Tw}_{\kappa}=+12$
Recall that it is conjectured that the space of domino tilings of a 3d box is connected via flips and trits. We now show that there is no fixed finite set of local moves under which the set of slab tilings of arbitrary boxes becomes connected. We first study the family of examples in Figures 20 and 21.
@@
Consider the region ${\cal R}=[0,2k]\times[0,2k]\times[0,5l]$, $k\geq 3$, $l\geq 1$ and a slab tiling ${\mathbf{t}}$ following the pattern in Figure 20 (piling copies if $l>1$). Now consider ${\cal R}_{0}=[0,2(k-1)]\times[0,2(k-1)]\times[0,5l]$. In ${\cal R}$, there is only one way of removing and placing back the slabs which are contained in the interior of ${\cal R}_{0}$.
@@
Refer to Figure 22. The region ${\cal R}_{0}\subset{\cal R}$ is bounded by the red curve. The region ${\cal R}_{1}=[0,2(k-2)]\times[0,2(k-2)]\times[0,5l]\subset{\cal R}_{0}$ is bounded by the dotted curve. We proceed by induction on $k$. We show the inductive process for five consecutive floors. The slab numbered $1$ is clearly stuck since it is “sandwiched” between two other slabs. Using this argument, on the first three floors, we proceed to deduce the positions of slabs which intersect the columns (in the floor diagram) between the dotted and red lines. In the figure, this corresponds to the slabs from $1$ to $8$. We repeat this, but now for slabs intersecting the rows between the dotted and red lines. This corresponds to slabs from $9$ to $14$.
@@
Figure 22: The tiling from Figure 21, with labels
@@
Corollary 6.4.
Let ${\cal R}$ and ${\mathbf{t}}$ be as in Lemma 6.3. Let $\tilde{\mathbf{t}}\neq{\mathbf{t}}$ be another slab tiling of ${\cal R}$. Then there are at least $32k-$ 3 slabs in ${\mathbf{t}}$ which are not in $\tilde{\mathbf{t}}$.
Proof.
Consider tilings ${\mathbf{t}}$ and $\tilde{\mathbf{t}}$ as in the statement. Let ${\cal X}\subseteq{\cal R}$ be the union of slabs in ${\mathbf{t}}$ which are not in $\tilde{\mathbf{t}}$. Alternatively, ${\cal X}$ is the union of slabs in $\tilde{\mathbf{t}}$ which are not in ${\mathbf{t}}$. Let $d$ be the number of slabs in ${\cal X}$. We need to prove that $d\geq 2k-3$.
We may assume without loss that ${\cal X}$ has connected interior. Indeed, otherwise let $\hat{\cal X}\subsetneqq{\cal X}$ be the closure of one of the connected components of the interior of ${\cal X}$. The region $\hat{\cal X}$ is tiled by ${\mathbf{t}}$ and $\tilde{\mathbf{t}}$ in two different ways. Let $\hat{\mathbf{t}}$ coincide with $\tilde{\mathbf{t}}$ in $\hat{\cal X}$ and with ${\mathbf{t}}$ elsewhere. The new tiling $\hat{\mathbf{t}}$ is a better example than $\tilde{\mathbf{t}}$ (i.e., with a smaller value of $d$).
As in Lemma 6.3 let ${\cal L}_{\operatorname{NE}}={\cal R}\smallsetminus{\cal R}_{0}\subset{\cal R}$ be the L shaped region to the NE direction. Define similar subsets ${\cal L}_{\operatorname{SE}}$, ${\cal L}_{\operatorname{SW}}$ and ${\cal L}_{\operatorname{NW}}$. By Lemma 6.3, we have ${\cal X}\cap{\cal L}_{\operatorname{NE}}\neq\varnothing$; the intersections of ${\cal X}$ with the other three L shaped sets are likewise nonempty. Let ${\cal M}_{\operatorname{North}}={\cal L}_{\operatorname{NE}}\cap{\cal L}_{\operatorname{NW}}$; define similarly ${\cal M}_{\operatorname{South}}$, ${\cal M}_{\operatorname{East}}$ and ${\cal M}_{\operatorname{West}}$. We claim that at least one of the two conditions below hold:
${\cal X}\cap{\cal M}_{\operatorname{North}}\neq\varnothing\neq{\cal X}\cap{\cal M}_{\operatorname{South}},\qquad{\cal X}\cap{\cal M}_{\operatorname{East}}\neq\varnothing\neq{\cal X}\cap{\cal M}_{\operatorname{West}}.$
Indeed, assume for instance that ${\cal X}\cap{\cal M}_{\operatorname{West}}=\varnothing$: the conditions ${\cal X}\cap{\cal L}_{NW}\neq\varnothing$ and ${\cal X}\cap{\cal L}_{SW}\neq\varnothing$ imply ${\cal X}\cap{\cal M}_{\operatorname{North}}\neq\varnothing\neq{\cal X}\cap{\cal M}_{\operatorname{South}}$; the other cases are similar.
Assume without loss of generality that ${\cal X}\cap{\cal M}_{\operatorname{North}}\neq\varnothing\neq{\cal X}\cap{\cal M}_{\operatorname{South}}$. Consider the planes $p_{x}=\{x\}\times[0,2k]\times[0,5l]$, $22\leq x\leq 2k-$ 2, which appear as horizontal lines in Figures 20 and 21. Given $x$, search for the slabs in ${\mathbf{t}}|_{{\cal X}}$ or $\tilde{\mathbf{t}}|_{{\cal X}}$ which cross $p_{x}$. We claim there are at least two such slabs. Notice first that if there are zero such slabs then ${\cal X}$ becomes essentially disconnected and we have a contradiction as above. Suppose then that there exist exactly one such slab. The region ${\cal X}_{\operatorname{North}}={\cal X}\cap([0,x]\times[0,2k]\times[0,5l])$ is tiled (by slabs, by either ${\mathbf{t}}$ or $\tilde{\mathbf{t}}$): the volume $V$ of ${\cal X}_{\operatorname{North}}$ is therefore a multiple of $4$. The region minus two unit cubes is also tiled (by either $\tilde{\mathbf{t}}$ or ${\mathbf{t}}$) and therefore $V-2$ is also a multiple of $4$, a contradiction. This completes the proof of the claim.
The total number $2d$ of slabs of both ${\mathbf{t}}$ and $\tilde{\mathbf{t}}$ in ${\cal X}$ is therefore at least $2(2k-3)$, completing the proof. ∎
Remark 6.5.
The corollary above is not sharp, the best estimate appears to be $12k-$ 1 (for all $l\geq 1$). The above formulation is sufficient for our main objective: no fixed set of local moves (involving a bounded number of slabs) connects all slab tilings of large boxes. More than that: given a fixed set of local moves, for sufficiently large $k$ the tiling ${\mathbf{t}}$ above admits none of these moves. $\diamond$
7 Possible values of the triple twist
Lemma 7.1.
There exist constants $C_{+}>0$ and $C_{1}>0$ such that the following assertions hold: If ${\cal R}=[0,N]^{3}$ is a cube with even side lengths then $|\operatorname{TTw}({\mathbf{t}})|<C_{+}N^{4}$ for any slab tiling ${\mathbf{t}}$ of ${\cal R}$. Also for ${\cal R}=[0,N]^{3}$ we have $\lvert\operatorname{TTw}[{\cal R}]\rvert\leq C_{1}N^{12}$.
Proof.
Given a pair of colors $\kappa$, we bound $\lvert\operatorname{Tw}_{\kappa}[{\cal R}]\rvert$. The transformed region $\tilde{{\cal R}}$ is contained in a cube of side $N$. See Figures 18 and 23 for an example of the transformed region for $N=4$. In Figure 23, the floor is bounded by the dotted $4\times 4$ square. We show $3$ dominoes on the central column. In this example, $d_{y}=d_{z}=1$ and there is one effect.
$y$
$x$
Figure 23: A typical floor after transforming a $4\times 4\times N$ box
Fix $f$ a floor of $\tilde{{\cal R}}$, fix a column $c$ of $f$, and use axes as in Figure 23. Consider dominoes with a square in $c$: there are $d_{y}$ dominoes in the direction of the $y$-axis and $d_{z}$ dominoes in the direction of the $z$-axis. The number of possible effects happening within $c$ is at most $d_{y}d_{z}$. Since $d_{y}+d_{z}\leq N$, we have $d_{y}d_{z}\leq\frac{N^{2}}{4}$ by AM-GM inequality.
Now, varying $c$ and then $f$, we see there are at most $\frac{N^{4}}{4}$ effects within $\tilde{{\cal R}}$. Thus, for every tiling ${\mathbf{t}}$ we have $|\operatorname{Tw}_{\kappa}({\mathbf{t}})|\leq\frac{N^{4}}{16}$. The first claim therefore holds for $C_{+}=\frac{1}{4}$.
Since $N\geq 2$, we have
This completes the proof, with $C_{1}=\frac{1}{2^{8}}$. ∎
Lemma 7.2.
Proof.
Consider the subdivision of the $[0,6n]\times[0,6n]\times[0,8n]$ box in parts $\alpha,\beta$ and $\gamma$ as in Figure 24. We tile $\alpha$ with horizontal slabs. We can tile $\beta$ and $\gamma$ both as collections of $2n^{2}$ annuli. In Figure 24, there are four kinds of floors. Each of them appears $2n$ times. Each rectangle on each floor has dimensions either $6n\times 2n$, $2n\times 6n$ or $2n\times 2n$.
$y$
$x$
$\alpha$
$\alpha$
$\beta$
$\alpha$
$\alpha$
$\alpha$
$\alpha$
$\beta$
$\beta$
$\gamma$
$\alpha$
$\alpha$
$\alpha$
$\alpha$
$\alpha$
$\gamma$
$\gamma$
$\beta$
$\alpha$
$\alpha$
$\gamma$
Figure 24: Floor diagram of subdivision of the $[0,6n]\times[0,6n]\times[0,8n]$ box
Consider the surfaces $S_{\beta}$ and $S_{\gamma}$ which are both squares of side $2n$, depicted as red cuts in Figure 25. Note that the top left vertex of a floor on the upper (resp. lower) diagram is of the form $(0,*,0)$ (resp. $(*,0,0)$). The surfaces $S_{\beta}$ and $S_{\gamma}$ are depicted in red.
$S_{\gamma}$
$z$
$x$
$S_{\beta}$
$y$
$z$
Figure 25: Example of $\beta$ ($y$-axis view) and $\gamma$ ($x$-axis view) on the $[0,12]\times[0,12]\times[0,16]$ box
Look at $S_{\beta}$ from the $y$-axis view. We regard each floor of $S_{\beta}$ with only grey pieces (a grey floor) as a domino tiling of an annulus, and we consider the flux with respect to the red cut. The flux across $S_{\beta}$ is the sum of the fluxes on all grey floors. The flux across $S_{\gamma}$ is defined analogously. We see the fluxes across each of the surfaces $S_{\beta}$ and $S_{\gamma}$ are integers in the interval $[-n^{2},n^{2}]$. We can choose the fluxes across $S_{\beta}$ and $S_{\gamma}$ to be equal to $u$ and $v$, respectively. We call this tiling ${\mathbf{t}}={\mathbf{t}}_{u,v}$.
We now prove $\operatorname{Tw}_{\kappa_{z}}({\mathbf{t}})=\pm 2uv$. We will consider $\tilde{{\mathbf{t}}}$ as a dimer covering ${\mathbf{t}}^{*}$ on the graph ${\cal R}^{*}$ whose vertices are centers of squares in $\tilde{{\cal R}}$ and edges join centers of adjacent squares. We draw ${\cal R}^{*}$ on top of ${\cal R}$ for ease of visualization. We put red dots on the squares of ${\cal R}$ painted with good colors. We proceed to define surfaces $\tilde{S}_{\beta}$ and $\tilde{S}_{\gamma}$ as follows. From the $z$-axis view, look at the first floor of $\beta$ which intersects $S_{\beta}$. Consider the red dots immediately above or below the red cut and draw a polygonal line joining them from left to right as in Figure 26. Then, $\tilde{S}_{\beta}$ is a curtain with base this polygonal line and height $12n-$ 1. The definition of the surface $\tilde{S}_{\gamma}$ is analogous: we now consider red dots to the left or right of the red cut, also in Figure 26. In Figure 26, the polygonal lines of $\tilde{S}_{\beta}$ and $\tilde{S}_{\gamma}$ are dotted and the surfaces $S_{\beta}$ and $S_{\gamma}$ are represented as full lines.
$x$
$y$
Figure 26: Construction of the surfaces $\tilde{S}_{\beta}$ (above) and $\tilde{S}_{\gamma}$ (below) following the example in Figure 25
To prove this, without loss, let ${\mathbf{t}}_{0}$ be ${\mathbf{t}}$ after twisting one of the annuli of $\beta$. The difference $\tilde{\mathbf{t}}_{0}-\tilde{{\mathbf{t}}}$ has one non trivial cycle which bounds an oriented surface $\tilde{S}$. We draw $\tilde{\mathbf{t}}_{0}-\tilde{{\mathbf{t}}}$ on top of ${\cal R}$ as in Figure 27. The surface $\tilde{S}$ bounded by $\tilde{\mathbf{t}}_{0}-\tilde{{\mathbf{t}}}$ is drawn as black polygonal lines. Full lines represent the first and last floors of this cycle. The first and last floors of two annuli from $\tilde{\gamma}$ are shown as purple polygonal lines. Full segments represent the actual dimers in the tiling. Notice $\tilde{S}$ is tangent to $\tilde{S}_{\gamma}$. To calculate $t=\operatorname{Tw}(\tilde{\mathbf{t}}_{0})-\operatorname{Tw}(\tilde{{\mathbf{t}}})$, we must count dominoes from $\tilde{{\mathbf{t}}}$ passing through $S$. These are exactly the ones which are part of an annulus in $\tilde{\gamma}$. By twisting one annulus from $\tilde{\gamma}$ we see $t$ changes by ±$\pm 2$, hence $t=\pm 2v$ by induction.
Figure 27: Some floors of the $[0,12]\times[0,12]\times[0,16]$ box
$S_{\gamma}$
$S_{\beta}$
Figure 28: We continue with the example of Figure 25
Without loss of generality we treat the first case of the above claim. Refer to Figure 28: on the first row we make the slabs in $\beta$ horizontal, on the second row we make the slabs in $\gamma$ and its hole horizontal. If this holds, then we can flip the slabs in $\beta$ so that all of them are horizontal. In particular, the slabs passing through $\gamma$’s hole are horizontal. Look at $\gamma$ from the $x$-axis view. For every pair of consecutive floors of $\gamma$ we can make all slabs horizontal by using the domino flip connectedness result for planar regions [5, 8]. ∎
Lemma 7.3.
Let ${\cal R}=[0,16n]^{3}$ be a box. For any triple $t\in([-2n^{4},2n^{4}]\cap 2{\mathbb{Z}})^{3}$ we construct a slab tiling ${\mathbf{t}}$ of ${\cal R}$ such that $\operatorname{TTw}({\mathbf{t}})=t$.
Proof.
For numbers $u_{i,j},v_{i,j}\in[-n^{2},n^{2}]$, $1\leq i\leq 2$, $1\leq j\leq 3$ we can construct tilings ${\mathbf{t}}_{i,j}$ of the $[0,6n]\times[0,6n]\times[0,8n]$ box such that $\operatorname{TTw}({\mathbf{t}}_{j})=2u_{i,j}v_{i,j}\mathbf{e_{j}}$ by Lemma 7.2 (here $\mathbf{e_{j}}\in{\mathbb{Z}}^{3}$ is the $j$-th vector in the canonical base). After doing this, we divide the $[0,16n]^{3}$ box in $8$ cubes of side $8n$. Each of these smaller cubes is tiled using one of the ${\mathbf{t}}_{i,j}$ on one corner, and horizontal slabs to fill the rest of the space. Two of the smaller cubes will receive only horizontal slabs. This new tiling is named ${\mathbf{t}}$.
Let $t=(t_{1},t_{2},t_{3})$. Choose
Here, for $s\in{\mathbb{R}}$, we have $s=\lfloor s\rfloor+\{s\}$ where $\lfloor s\rfloor\in{\mathbb{Z}}$ and $\{s\}\in[0,1)$. We have $t_{j}=2u_{1,j}v_{1,j}+2u_{2,j}v_{2,j}$, as desired. Then, for each $j$, $\operatorname{TTw}({\mathbf{t}}_{1,j}+{\mathbf{t}}_{2,j})=t_{j}\mathbf{e_{j}}\in([-2n^{4},2n^{4}]\cap 2{\mathbb{Z}})^{3}$, where ${\mathbf{t}}_{1,j}+{\mathbf{t}}_{2,j}$ is the tiling of the two cubes containing ${\mathbf{t}}_{1,j}$ and ${\mathbf{t}}_{2,j}$. Summing the above for all $1\leq j\leq 3$ yields $\operatorname{TTw}({\mathbf{t}})=(t_{1},t_{2},t_{3})$. ∎
Lemma 7.4.
Proof.
Write $N=16M+2k$ for some $0\leq k\leq 7$. We can divide ${\cal R}$ into two parts ${\cal R}_{1}$ and ${\cal R}_{2}$ as in Figure 29 (there is one kind of floor: it comprises one $16M\times 16M$ square on the lower left corner, ${\cal R}_{1}$, and its complement, ${\cal R}_{2}$). For a given $\mathbf{v}\in([-2M^{4},2M^{4}]\cap 2{\mathbb{Z}})^{3}$, we tile ${\cal R}_{1}$ with a tiling ${\mathbf{t}}_{1}$ having $\operatorname{TTw}({\mathbf{t}}_{1})=\mathbf{v}$ and we tile ${\cal R}_{2}$ with horizontal slabs.
$\mathcal{R}_{1}$
$\mathcal{R}_{2}$
Figure 29: Diagram of subdivision of ${\cal R}$
Then, $\lvert\operatorname{TTw}[{\cal R}]\rvert\geq(2M^{4})^{3}=8M^{12}=8(\frac{N-2k}{16})^{12}=c(N-2k)^{12}$, where $c=\frac{1}{2^{45}}$. We seek $C>0$ such that $(N-2k)^{12}\geq CN^{12}$ holds. This happens if and only if $C\leq(1-\frac{2k}{N})^{12}$. By minimizing the RHS with the constraints $N\geq 256$ and $k\leq 7$, we can take $C=\frac{1}{2}\leq(1-\frac{14}{256})^{12}$. Finally, we ca ${\cal R}=[0,N]^{3}$n take $C_{0}=cC=\frac{1}{2^{46}}$. ∎
We summarize the previous lemmas in the proposition below.
Proposition 7.5.
If $N\geq 256$ is an even integer, then
Proof.
We take $C_{0}$ and $C_{1}$ as in the proofs of Lemmas 7.1 and 7.4. ∎
We are also ready for Theorem 1.
Proof of Theorem 1.
We use the max norm and give explicit values for $C_{\pm}$ assuming $N\geq 256$. As in Lemma 7.1, we can take $C_{+}=\frac{1}{4}$. We claim that $C_{-}=1/{2^{16}}$ works. Indeed, take $n=\left\lfloor\frac{N}{16}\right\rfloor$ and consider the smaller cube ${\cal R}_{-}=[0,16n]^{3}$. Notice that $N^{4}<2^{17}n^{4}$. Each coordinate of $t$ is an even integer with absolute value smaller than $C_{-}N^{4}<2n^{4}$. Apply Lemma 7.3 to deduce that there exists a tiling ${\mathbf{t}}_{-}$ of ${\cal R}_{-}$ with $\operatorname{TTw}({\mathbf{t}}_{-})=t$. Finally, construct a tiling ${\mathbf{t}}$ of ${\cal R}=[0,N]^{3}\supseteq[0,16n]^{3}$ by starting with ${\mathbf{t}}_{-}$ and completing the remaining region with horizontal slabs. We have $\operatorname{TTw}({\mathbf{t}})=\operatorname{TTw}({\mathbf{t}}_{-})=t$, as desired. ∎